Solve the following equations for equilibrium price and quantity.

Demand Curve: Qdx = 10 -2Px

Supply Curve: Qsx = 5 + 3Px
Answer #1

Answer

the equilibrium is at Qd=Qs

10-2P=5+2P

5P=5

P=1

Q=10-2*1=8 units

the equilibrium price is $1 and the quantity is 8 units